# Lightweight LLM Agent Memory with Small Language Models

### 09 Apr 2026 v1

@@ Abstract
Jiaquan Zhang1, Chaoning Zhang1,*, Shuxu Chen2, Zhenzhen Huang1, Pengcheng Zheng1, Zhicheng Wang1, Ping Guo4, Fan Mo5, Sung-Ho Bae3, Jie Zou1, Jiwei Wei1, Yang Yang1 1School of Computer Science and Engineering, University of Electronic Science and Technology of China; 2Department of Electronics and Information Convergence Engineering, Kyung Hee University; 3School of Computing, Kyung Hee University; 4Computer Science Department, City University of Hong Kong; 5Department of Computer Science, University of Oxford *Correspondence: chaoningzhang1990@gmail.com
@@ 1 Introduction
LLM-driven agents excel at long-term dialogue, multi-step reasoning, and task-oriented interaction Zhang et al. (2026c); Huang et al. (2024); Zhu et al. (2025); Zhang et al. (2026a, b, 2025a). To maintain cross-turn consistency beyond the context window, many systems augment agents with external memory Lee et al. (2024); Xu et al. (2025); Hu et al. (2025); Wang et al. (2026). Long-term memory supports continual learning, personalization, and planning.
@@
To address the above questions, we propose LightMem, a lightweight memory system for LLM agents that models long-term memory as an incrementally evolvable process via specialized SLMs. LightMem modularizes and decouples query parsing, memory retrieval, memory writing, and long-term consolidation, enabling independent optimization and elastic scaling of each component. This separation allows for lightweight online processing to be distinct from offline consolidation, achieving efficient memory use under tight compute budgets and supporting long-term evolution. Memories are organized into STM, MTM, and LTM stores based on temporal and access characteristics. User-identity metadata is embedded in each memory unit to enforce user-level logical isolation, balancing privacy, consistency, and scalability. We propose a modular online–offline memory pipeline in which distinct SLMs specialize in complementary memory operations. Online, LightMem uses three specialized small language model modules: a Controller (SLM-1) for intent and query planning, a Selector (SLM-2) for candidate verification and compression, and a Writer (SLM-3) for incremental memory writing. Specifically, the Controller rewrites the user input into intent-conditioned hypothetical queries (HQs) and allocates a fixed Top-$K$ budget across MTM and LTM. Given the retrieved candidates, the Selector performs metadata-constrained prefiltering followed by semantic-consistency based re-ranking to output the final Top-$K$ memories. After each turn, the Writer summarizes the interaction into compact MTM entries and maintains MTM incrementally. Long-term abstraction and consolidation are handled offline by a large-context model, keeping the online path lightweight. Our main contributions are:
@@ 3.3 Intent Modeling and Retrieval Control
Given $(x_{t},C_{t})$, SLM-1 serves as a lightweight retrieval controller rather than a retriever. It converts the raw input into a structured retrieval plan that specifies: (i) what to search (queries), (ii) how to constrain the search (metadata filters), and (iii) how many items to return under a fixed Top-$K$ budget. Concretely, SLM-1 first infers coarse intent attributes, e.g., whether the query depends more on recent episodic details or long-term stable knowledge, and whether strong personalization is needed. These attributes are used only for retrieval planning (query rewriting, routing, and budgeting), not for answer generation. SLM-1 then rewrites the input into a set of hypothetical queries $\{q_{t}^{(i)}\}$ and outputs metadata constraints $\phi_{t}$ (user identifier, optional time window, and type tags) to reduce noise and enforce user-level isolation. Finally, it issues the retrieval request:
@@ 4.1.3 Metric
We evaluate response quality using both lexical-overlap and semantic metrics. Specifically, we report F1 and BLEU-1 to measure answer correctness and token-level overlap with reference responses, which are commonly used for short factual and span-like outputs. For DialSim Kim et al. (2024), where multiple surface forms can express the same meaning, we additionally report ROUGE-L and METEOR to capture sequence-level overlap and soft token matching, and use SBERT similarity to quantify semantic consistency between generated and reference responses beyond n-gram matches. These metrics offer a dependable evaluation of both precision and meaning retention within long-horizon dialogue scenarios.
@@ 4.1.4 Implement Details
Generate HQs Rewrite the user request into one or more standalone hypothetical queries, optionally splitting intent into user-specific vs. factual queries, and making implied context (time/location) explicit.
@@ 4.2 Performance
Performance. As shown in Table 2, LightMem achieves the best overall performance on the LoCoMo benchmark. In comparison to baselines that directly rely on long-context replay (such as LoCoMo and MemGPT), LightMem demonstrates more stable performance across the majority of question categories. It has particularly distinct advantages in multi-hop and temporal questions, all while not being reliant on a single long-context input. In comparison with summarization- or compression-based methods (MemoryBank and ReadAgent), LightMem consistently outperforms them across all categories, with larger margins on multi-hop and temporal reasoning. When compared to the strongest memory-centric baseline, A-MEM, LightMem further improves overall performance in most model settings. On GPT-4o, LightMem attains an F1 score of 34.50 for multi-hop questions, surpassing A-MEM’s score of 32.86. Overall, LightMem demonstrates more robust long-term memory across different model scales and question types. In addition, it maintains significant advantages in cross-session reasoning tasks.
@@ 4.3 Ablation Study
Experimental setup. Under a unified deployment configuration, we assess the scalability and latency by utilizing GPT-4o-mini. Subsequently, we present the following latency metrics:
@@ 5 Conclusion
We present LightMem, a lightweight external memory system for LLM agents driven by small language models (SLMs). LightMem decouples high-frequency online memory operations (query control, retrieval, and writing) from offline consolidation, and organizes memory into STM/MTM/LTM with user-scoped isolation. Online, it operates under a fixed retrieval budget and uses HQ-based routing plus a two-stage retrieval (coarse vector search followed by semantic consistency re-ranking) to reduce retrieval noise. Experiments on LoCoMo and DialSim show consistent gains across model scales, including  +2.5 average F1 on LoCoMo, improved semantic consistency on DialSim, and low median latency (83 ms retrieval; 581 ms end-to-end).

### 19 Apr 2026 v2

@@ Abstract
Jiaquan Zhang1, Chaoning Zhang1,*, Shuxu Chen2, Zhenzhen Huang1, Pengcheng Zheng1, Zhicheng Wang1, Ping Guo3, Fan Mo4, Sung-Ho Bae2, Jie Zou1, Jiwei Wei1, Yang Yang1 1University of Electronic Science and Technology of China; 2Kyung Hee University; 3City University of Hong Kong; 4University of Oxford *Corresponding author: chaoningzhang1990@gmail.com
@@ 1 Introduction
LLM-driven agents excel at long-term dialogue, multi-step reasoning, and task-oriented interaction Zhang et al. (2026c); Huang et al. (2024); Zhu et al. (2025); Zhang et al. (2026a, b, 2025a). To maintain cross-turn consistency beyond the context window, many systems augment agents with external memory Lee et al. (2024); Xu et al. (2025); Hu et al. (2025); Wang et al. (2026). Long-term memory supports continual learning and planning.
@@
To address the above questions, we propose LightMem, a lightweight memory system for LLM agents that models long-term memory as an incrementally evolvable process via specialized SLMs. LightMem modularizes and decouples query parsing, memory retrieval, memory writing, and long-term consolidation, enabling independent optimization and elastic scaling of each component. This separation allows for lightweight online processing to be distinct from offline consolidation, achieving efficient memory use under tight compute budgets and supporting long-term evolution. Memories are organized into STM, MTM, and LTM stores based on temporal and access characteristics. User-identity metadata is embedded in each memory unit to enforce user-level logical isolation, balancing privacy, consistency, and scalability. We propose a modular online–offline memory pipeline in which distinct SLMs specialize in complementary memory operations. Online, LightMem uses three specialized small language model modules: a Controller (SLM-1) for intent and query planning, a Selector (SLM-2) for candidate verification and compression, and a Writer (SLM-3) for incremental memory writing. Specifically, the Controller rewrites the user input into intent-conditioned hypothetical queries (HQs) and allocates a fixed Top-$K$ budget across MTM and LTM. Given the retrieved candidates, the Selector performs metadata-constrained prefiltering followed by semantic-consistency based re-ranking to output the final Top-$K$ memories. After each turn, the Writer summarizes the interaction into compact MTM entries and maintains MTM incrementally. Long-term abstraction and consolidation are handled offline by a large-context model, keeping the online path lightweight. Our main contributions are summarized as:
@@ 3.3 Intent Modeling and Retrieval Control
Given $(x_{t},C_{t})$, SLM-1 serves as a lightweight retrieval controller rather than a retriever. It converts the raw input into a structured retrieval plan that specifies: (i) what to search (queries), (ii) how to constrain the search (metadata filters), and (iii) how many items to return under a fixed Top-$K$ budget. Concretely, SLM-1 first infers coarse intent attributes, e.g., whether the query depends more on recent episodic details or long-term stable knowledge, and whether strong personalization is needed. These attributes are used only for retrieval planning (query rewriting, routing, and budgeting), not for answer generation. SLM-1 then rewrites the input into a set of HQs $\{q_{t}^{(i)}\}$ and outputs metadata constraints $\phi_{t}$ (user identifier, optional time window, and type tags) to reduce noise and enforce user-level isolation. Finally, it issues the retrieval request:
@@ 4.1.3 Metric
We evaluate response quality using both lexical-overlap and semantic metrics. Specifically, we report F1 and BLEU-1 to measure answer correctness and token-level overlap with reference responses, which are commonly used for short factual and span-like outputs. For DialSim Kim et al. (2024), where multiple surface forms can express the same meaning, we additionally report ROUGE-L and METEOR to capture sequence-level overlap and soft token matching, and use SBERT similarity to quantify semantic consistency between generated and reference responses beyond n-gram matches. These metrics provide a reliable assessment of both precision and meaning retention in long-horizon dialogue scenarios.
@@ 4.1.4 Implement Details
Generate HQs Rewrite the user request into one or more standalone HQs, optionally splitting intent into user-specific vs. factual queries, and making implied context (time/location) explicit.
@@ 4.2 Performance
Performance. As shown in Table 2, LightMem achieves the best overall performance on the LoCoMo benchmark. In comparison to baselines that directly rely on long-context replay (such as LoCoMo and MemGPT), LightMem demonstrates more stable performance across the majority of question categories. It has particularly distinct advantages for multi-hop and temporal questions, all while not relying on a single long-context input. In comparison with summarization- or compression-based methods (MemoryBank and ReadAgent), LightMem consistently outperforms them across all categories, with larger margins on multi-hop and temporal reasoning. When compared to the strongest memory-centric baseline, A-MEM, LightMem further improves overall performance in most model settings. On GPT-4o, LightMem attains an F1 score of 34.50 for multi-hop questions, surpassing A-MEM’s score of 32.86. Overall, LightMem demonstrates more robust long-term memory across different model scales and question types. In addition, it maintains significant advantages in cross-session reasoning tasks.
@@ 4.3 Ablation Study
4.4 Scalability and Latency
Under a unified deployment configuration, we assess the scalability and latency by utilizing GPT-4o-mini. Subsequently, we present the following latency metrics:
@@ 4.4 Scalability and Latency
4.5 Performance Stability under MTM Growth
We analyze how MTM size affects performance stability on DialSim with Llama-3.2-1B. Rather than rerunning the model with different memory sizes, we report cumulative F1 at several natural growth stages along the same full dialogue trajectory, approximately at 100, 1,000, 5,000, and 10,000 MTM entries. As shown in Table 5, the gap between the two methods is small when MTM remains relatively small. However, as memory entries accumulate along the dialogue trajectory, pure vector retrieval becomes increasingly affected by retrieval noise, whereas LightMem maintains more stable performance. This trend is consistent with the role of Stage 2 semantic filtering, which helps preserve retrieval quality as the memory store grows.
Table 5: Performance stability under natural MTM growth on DialSim with Llama-3.2-1B. Statistics are computed cumulatively along the same full dialogue trajectory at different MTM sizes.
[…]
$\sim$100 3.95 3.98 0.03
$\sim$1,000 3.90 4.05 0.15
$\sim$5,000 3.86 4.09 0.23
10,000 3.83 4.12 0.29
4.6 Error Injection Stress Test
To study error accumulation in long-horizon memory systems, we conduct an end-to-end error injection stress test on DialSim under a fixed final Top-$K$ memory budget. Group A is the full LightMem system. Group B injects 50% irrelevant HQs to simulate SLM-1 failures in query planning. Group C removes SLM-2 semantic reranking and directly uses the stage-one vector Top-$K$. Group D replaces 50% of SLM-3 writes with noisy strings, simulating persistent MTM pollution. Group E combines all three perturbations to simulate cascading failure. As shown in Table 6, performance degrades consistently as errors are introduced into different stages of the pipeline. Query noise in SLM-1 causes a moderate drop, suggesting partial robustness to retrieval-side perturbations. Removing SLM-2 leads to a larger decline, confirming the importance of semantic filtering after coarse retrieval. Noisy writes in SLM-3 further degrade performance, indicating that memory pollution introduces persistent downstream effects by contaminating future retrieval. Cascading failure causes the largest collapse, showing that errors can accumulate and amplify across turns when noisy retrieval, missing filtering, and corrupted writing occur simultaneously.
Table 6: Error injection stress test on DialSim under a fixed final Top-$K$ memory budget.
Group Setting F1 SBERT
A Full LightMem 4.12 23.40
B SLM-1 50% HQ noise 3.95 23.08
C w/o SLM-2 reranking 3.83 22.56
D SLM-3 50% write noise 3.78 22.45
E Cascading failure 1.85 11.20
4.7 Update Gap Stress Test
As shown in Table 7, concurrent retrieval from MTM and LTM is important for handling the update gap. The full LightMem system achieves the best performance, with a multi-hop F1 of 28.85, whereas using only LTM or only MTM results in clear performance drops. This suggests that concurrent routing effectively bridges recent unconsolidated facts in MTM and older consolidated knowledge in LTM. A more challenging case appears in Group D. Under MTM noise saturation, the multi-hop F1 decreases to 23.10, indicating that heavy recent noise can interfere with retrieval during the update gap. This is likely because the fixed Top-$K$ budget is increasingly occupied by irrelevant recent records, making the newest relevant fact harder to retrieve before it is consolidated into LTM. Overall, the results show that concurrent retrieval provides a practical and effective solution for update-gap reasoning, especially when recent and long-term evidence must be combined.
Table 7: Update gap stress test on the LoCoMo multi-hop subset with GPT-4o-mini.
Group Setting Multi-hop F1
A Full LightMem (normal gap) 28.85
B LTM-only retrieval 19.45
C MTM-only retrieval 20.12
D MTM noise saturation 23.10
@@ 5 Conclusion
We present LightMem, a lightweight external memory system for LLM agents driven by SLMs. LightMem decouples high-frequency online memory operations (query control, retrieval, and writing) from offline consolidation, and organizes memory into STM/MTM/LTM with user-scoped isolation. Online, it operates under a fixed retrieval budget and uses HQ-based routing plus a two-stage retrieval (coarse vector search followed by semantic consistency re-ranking) to reduce retrieval noise. Experiments on LoCoMo and DialSim show consistent gains across model scales, including  +2.5 average F1 on LoCoMo, improved semantic consistency on DialSim, and low median latency (83 ms retrieval; 581 ms end-to-end).
@@ 6 Limitations
7 Acknowledgements
This work was partially supported by the National Natural Science Foundation of China under grant 62572104, and 62220106008.